Fermi National Laboratory
 

Accelerator Update for June 3, 2009 - June 5, 2009

 

Wednesday June 3

The day shift began with Operations monitoring the Tevatron (TeV) store 7102 (34.23E30) and the Recycler stash (144.4E10), with the Pelletron down for repairs, with the Antiproton Source stacking (145.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

Pelletron personnel completed their high voltage repairs and had the system online by 11 AM.

Operations terminated store 7102 at 11:30 AM. TeV personnel conducted a one-hour end-of-store study.

Operations sent beam to the next Meson MTest experiment (T988) when beam was available.

The TeV study was followed by a TeV sector E1 wet engine flywheel cryo system repair, which was completed by 5 PM.

Operations checked the TeV orbits and then put the TeV into HEP shot setup at 7:15 PM.

Operations established store 7105 at 8:30 PM with an average initial luminosity of 284E30.

Operators tuned Booster and MI.

Pbar transferred antiprotons to the Recycler when appropriate.

Thursday June 4

The midnight shift began with Operations monitoring store 7105 (160.35E30) and the stash (271.0E10), with Pbar stacking (20.9E10), and with NuMI and MiniBooNE taking beam.

Operations sent beam to Meson MTest when beam was available.

Operations terminated store 7105 at 9:30 AM and then put the TeV into HEP shot setup.

Operations established store 7106 at 11:45 AM with an average initial luminosity of 279E30.

Operations halted the Booster and Preaccelerator beam for around 30 minutes so Booster personnel could install a power supply cable and Preaccelerator personnel could access the H- Source dome.

Store 7106 aborted when the TeV sector A3 VCB opened up. There was no quench. EE Support personnel swapped out the safety coordinator module at sector A3 with the one at B2.

The TeV recovered from the abort and Operations started shot setup at 7:40 PM.

Operations established store 7109 at 9 PM with an average initial luminosity of 266E30.

Pbar transferred antiprotons to the Recycler when appropriate.

Friday June 5

The midnight shift began with Operations monitoring store 7109 (158.71E30) and the stash (91.9E10), with Pbar stacking (26.7E10), and with NuMI and MiniBooNE taking beam.

Operations terminated store 7109 at 5:20 AM after a TeV sector A4 house air compressor tripped off. Shortly after the store was terminated the sectors A4 and B0 vacuum valves closed due to a lack of air pressure. Operations called in Mechanical Support personnel to install a back up air compressor.

After the air compressor trip, the output of a TeV quadrupole power supply (D0Q2) jumped and caused a sector DA cryo system quench. Operations called in EE Support personnel to investigate the power supply. Efforts to understand why the power supply output changed continued through out the midnight shift.

At the end of the midnight shift, a TeV A-E sector access was being planned to investigate why the A4 cryo and vacuum valves would not open remotely.

Pbar transferred antiprotons to the Recycler when appropriate.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow some necessary studies in the TeV and MI today.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/05/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory